Calculate the molarity of a solution of HCl when 87 g of HCl are mixed with 230 mL of
H2O.
1
Expert's answer
2020-05-15T10:05:44-0400

m(HCl) = 87 g

M(HCl) = 36.5 g/mol

V(H2O) = 230 mL = 0.23 L

ρ(HCl) = 1.48 g/L

n(HCl) = m(HCl)/M(HCl) = 87 g/ 36.5 g/mol = 2.38 mol

V(solution) = m(HCl)/ρ(HCl) + V(H2O) = 58.78 + 0.23 = 59.01 L

C(solution) = n(HCl)/V(solution) =

= 2.38 mol/ 59.01 L = 0.04 M